# HCT-QA: A Benchmark for Question Answering on Human-Centric Tables

### v2

@@ 3.1 Real-world dataset collection
(c) Main properties of synthetic HCTs (d) Main properties of real-world HCTs
Figure 2: Main properties of the HCTs and QA pairs in our benchmark datasets.
3.2 Query and ground-truth answer generation
In our context, a query on an HCT object, is a natural language question to extract some information, either directly (e.g., selecting some values from the table), or after applying some computations (e.g., expression, aggregation). We focus on single-table queries. Our benchmark has two types of queries: (1) Human-Generated are curated by domain experts; (2)  Model-Generated are generated by an LLM. All query-answer pairs underwent a human verification to ensure high quality.
Characterization of query properties:
We developed a set of properties, inspired by typical SQL operations, that can be used to describe each query. These properties capture information about the number and nature of the operations, such as aggregation and filtering, that are required to answer the query. For instance, the number of filtered rows (WHERE clause) and projected columns (Condition), the number of returned columns (SELECT clause), or more complex properties involving aggregations and ranks (see Appendix B for more details). These properties are useful in understanding what sorts of queries models may struggle with the most. In our benchmark, properties are only provided for human-generated queries.
Human-Generated Queries:
To generate a query and its corresponding answer, the domain expert uses a custom annotation tool (Appendix B) to see the image of the table, create the query and answer, and annotate the properties of that query. We also allow the user to select among a set of pre-defined query templates to create a question that adheres to the query properties of that template and find its answer in the HCT, enabling automatic filling of the query properties. The expert may see the type of queries they have created so far as a nudge to create a more balanced set of queries that require various operators, expressions, and aggregations. Each query was verified by another domain expert for quality control. We ensure that each HCT has at least one human-generated query. We created 3,248 such queries on Real World HCTs and found that generating each query takes around 2.5 minutes.
Model-Generated Queries:
To expand our query set, we leverage the OpenAI’s GPT model  openai2023gpt4-turbo . The model is provided with the HCT image, an example human-generated query-answer pair (one-shot prompting), and instructions containing the constraints the query must adhere to. The model is then prompted to return a set of 8 query-answer pairs for each table. This method allows us to scale the number of queries very well. We generated 6,587 such queries on Real World HCTs in approximately 30 minutes.
Human Verification:
Model-generated queries were given to a team of annotators to: (i) rate the quality of the question as ”good”, ”ambiguous”, ”nonsensical” and (ii) given the table and query, correct the corresponding answer if it is wrong (check Appendix B for details). We dropped ”nonsensical” queries, rewrote ”ambiguous” queries and kept the ”good” ones. Each query was independently verified by two annotators. We included the human-generated queries as well to gauge the quality of the annotators and removed annotators that marked more than 1 manually generated queries as ”nonsensical” or its answer as incorrect. This process ensured that we have a set of high-quality query-answer pairs. We found that only 4% of model-generated queries were marked as ”nonsensical”, 24% as ambiguous, and 72% as ”good”. Moreover, the model required answer corrections for nearly 50% of the queries. Finally, we obtained $1,880$ tables and $9835$ related question and answer pairs from QNPC, Pakistan Census, US Census, and ARXIV real-world sources (Figure 2).
3.3 Synthetic tables and QA generation
To expand the dataset, we developed a synthetic data generation tool to produce HCTs in CSV, HTML, and PDF formats, along with their queries and answers, aw well as their properties. Synthesizing datasets in benchmarking is significantly useful in various ways including filling in gaps such as specific table patterns that do not sufficiently exist in real-world data, stress testing certain combinations, scaling up the data size, and finally by controlling the data generation (for both HCTs and their associated queries) the ground truth answers become computationally known without the need for manual validation.
We generated $4,679$ unique synthetic HCTs ($T_{HCT}$) across $7$ semantics, with up to $15$ template-based natural language questions $Q_{NL}$ per table, leading to $67,747$ triplets $(T_{HCT},Q_{NL},A)$ (Figure 2) with the corresponding ground-truth answer $A$. Each synthetic HCT contains no duplicate values to avoid ambiguities when evaluating the matching between the ground-truth and LLM-predicted answers. In practice, HCTs are produced by pivoting pivotTable2005 standard relational tables $T_{REL}$ that we first generated from custom semantic data templates. Attributes of $T_{REL}$ are randomly selected and ordered for the pivoting process (pivottabler, ) to create a variety of the HCT properties described in Section 3.1, such as forming nested rows and columns of $T_{HCT}$, as well as adding, naming, and positioning aggregate rows or columns. We automatically get the ground-truth answer $A$ to a SQL query $Q_{SQL}$ on the relational table $T_{REL}$ using a $SQL$ engine. We then transform the SQL query $Q_{SQL}$ into $Q_{NL}$ using custom templates to get the desired triplet $(T_{HCT},Q_{NL},A)$.
Our synthetic generator can be customized to produce 16 out of the 18 layout characteristics detailed in subsection 3.1, and the table semantics can be easily extensible to any domain. Further details are given in Appendix C.
4 Experimental evaluation
Our goal is to build a benchmark dataset on which one can test how natural language queries can be answered on HCTs using any data processing method or engine. In our experiments, we evaluate our benchmark on Large Language Models (LLMs) and Vision Language Models (VLMs). We want to answer the following questions: (RQ1) How do these models perform overall on HCT-QA? (RQ2) Are models improving over time? (RQ3) Does the input format of the HCT matter in text-modality? (RQ4) Does fine-tuning models for this task improve performance? (RQ5) How do HCT properties affect model performance? We also explored four additional questions (described in Appendix D.2): (RQ6) Does model verbosity affect the output correctness? (RQ7) How do question properties affect model performance? (RQ8) How are model size and performance correlated? (RQ9) In what cases do VLMs perform better than LLMs?
4.1 Experimental setup
Models.
@@ 4.1 Experimental setup
We tested our benchmark on a large variety of open- and closed-weight LLMs and VLMs. Details about these models are given in Table 8 (Appendix D.1).
Datasets.
Our benchmark comprises four real-world data sources (Real World HCTs) and one synthetically generated dataset (Synthetic HCTs) with their corresponding QA Pairs (Table 2). There are 1,880 real world HCTs with 9,835 QA pairs on them and 4,679 synthetic HCTs with 67,747 queries on them (total of 6,559 HCTs with 77,582 queries).
Prompting approach and hyperparameters.
@@
We experimented with zero-shot and one-shot prompting and found that there was no clear indication that the latter improved model performance. Thus, for all subsequent experiments, we used zero-shot prompting. Additional details on the prompting approaches and tests can be found in Appendix D.3. The temperature was set to 0, and the maximum output tokens to 128. We post-process the model responses using a rule-based approach to extract the actual answer if one was provided. All inference and fine-tuning experiments were run on 4 NVIDIA A100 GPUs.
Evaluation metrics.
We used three standard metrics to evaluate the performance of the model: (i) Precision, (ii) Recall, and (iii) Complete Containment (CC Score), which equals 1 if the model outputs the complete answer (Recall = 100%) and 0 otherwise. It is important to note that Precision may be unfairly deflated due to the verbose nature of language models. In the prompt, we specify outputting the answer in an easy-to-parse set-like format. However, although most models follow the specified set format, Precision would not capture the correctness of the answer if the model provides it as part of a sentence. Additionally, we checked the robustness of using an LLM-as-a-judge for evaluation and found it to closely align with the CC Score metric. We thus opted not to use it (see Appendix D.4). Our analysis focuses primarily on Recall scores to compare different models.
Table 1: Overall performance of models. All text-only and vision-text models are evaluated on the appropriate datasets. Confidence intervals are given in Appendix D.7
@@
Model Name QNPC PAK US CENSUS arXiv Synthetic Tables All Data
Precision Recall CC Score Precision Recall CC Score Precision Recall CC Score Precision Recall CC Score Precision Recall Precision Recall

### v3

@@ III-A Real-world HCTs
(c) Main properties of real-world QA pairs (d) Main properties of synthetic HCTs
Figure 3: Main properties of the HCTs and QA pairs in HCT-QA.
III-B Questions and Ground-Truth Answers Generation
A question over an HCT object refers to a NL question aimed at retrieving specific information—either directly (e.g., selecting particular value(s)) or through computation (e.g., expressions or aggregations). Our focus is on single-HCT questions. For the real-world dataset, we adopt two approaches for question generation: (1) Manually-generated questions – these are produced by the authors of this paper who are database experts with several years of combined experience in academia and industry, and (2) model-generated questions – these are produced by an LLM. All QA pairs, regardless of their source, are subsequently subject to independent human verification to maintain high dataset quality.
Characterization of Question Properties. We designed a set of properties, inspired by common SQL operations, to systematically characterize each question. These properties capture both the type and complexity of operations—such as filtering, aggregation, rank, and projection—required to derive the answer. For each operation, additional sub-properties are maintained to capture more details on the operation’s complexity. For example, the question properties include the number of filtering conditions, whether the projected columns include expressions or not, the number and type of aggregators (e.g., min, max, and avg.) as well as if they are local (a specific subset of values in the HCT) or global (values across the HCT), and the ranking on plain column or expression. These properties provide valuable insights into the types of questions that models tend to find most challenging. To facilitate capturing these properties, we designed a custom interface as depicted in Figure 4. These properties are only provided for the manually-generated questions.
Figure 4: Question interface for the properties of manually-generated questions
Manually-Generated QA Pairs. To generate a question and its answer, we use the annotation interface shown in Figure 4. This interface enables the user to visualize the table image, formulate a question with its associated answer, and annotate the relevant question properties. Also, the interface presents real-time statistics on the types of questions generated, encouraging the user to create a balanced distribution of questions involving diverse operators, expressions, and aggregations. Each question undergoes verification by two independent persons to ensure quality and consistency. We guarantee that every HCT includes at least one manually-generated question. In total, we produced 3,248 such questions with their properties based on real-world HCTs, with each question requiring approximately 1.5 to 3 minutes to complete.
Model-Generated QA Pairs. To expand our question set beyond, we use OpenAI’s GPT [46] to automatically generate additional questions. The model is provided with the HCT image, a sample manually-generated QA pair from the previous step (using one-shot prompting), and detailed instructions specifying the required constraints, such as the target operations and the desired level of question complexity. The model then produces a set of eight QA pairs per HCT, enabling efficient large-scale question generation. Using this approach, we generated 6,587 additional QA pairs on real-world HCTs in approximately 30 minutes.
To ensure the accuracy and clarity of model-generated questions, we established a rigorous human verification and correction pipeline. Each model-generated QA pair is reviewed by a team of annotators who (i) assess the quality of the question as good, ambiguous or nonsensical and (ii) correct the corresponding answer if it is incorrect or poorly formatted. Questions that are rated as nonsensical are excluded from the dataset, while those marked as ambiguous are re-written. Our annotation team comprised of 14 undergrad students majoring in Computer Science who had experience with databases. Annotators were given a training session by the authors on the task description and did supervised mock exercises of the task. Each QA pair is evaluated by two annotators independently. In cases of disagreement, a third annotator makes the final decision. To monitor the consistency and attentiveness of annotators, we embedded a set of pre-validated, manually-generated questions within their evaluation batches. Annotators who misclassified more than one of these verified questions (e.g., labeling them as ambiguous or nonsensical or incorrectly flagging their answers) were excluded from the annotation pool. Only a single annotator failed our hidden test questions and was replaced.
Our model-based pipeline for QA generation and verification was highly effective and scalable. Overall, only 4% of model-generated questions were marked as nonsensical, 24% as ambiguous, and 72% as good. Moreover, the model’s answers required minor corrections for nearly 50% of the questions. Combining the manually-generated and model-generated questions, we ended up with a total of 9,835 QA pairs over the 1,880 HCTs. In Figure 3(a), we present the distribution of the questions across the different datasets. Moreover, in Figure 3(b), we present aggregated distribution of the question properties over all questions. For example, roughly 2,000 questions involve aggregations, while 1,000 questions involve ranking, and there are 594 yes/no questions. Overall, Figure 3(b) shows a good distribution of NL questions on real-world HCTs ranging from simple yes/no questions to complex combination of aggregation, ranking, column expression, etc.
Manual Effort. Referring to Figure 2, to prepare 1,880 real-world HCTs, a combined 150 manual work hours (MWH) were spent as follows: 15 MWH for HCTs filtering to meet certain standards (e.g., visual quality and no out-of-scope properties), 15 MWH for sample verification (10% from text and image modalities), 60 MWH for annotating HCT properties, and 80 MWH for 2-way verification of the annotated HCTs and their properties. To prepare 9,835 QA pairs, 1020 MWH were spent as follows: 150 MWH for manually creating and annotating QA pairs and their properties, 270 MWH for 2-way verification of the manually created and annotated QA pairs, and 600 MWH to verify model generated QA pairs. The creation and verification of the manual QA pairs was done by the authors, whereas the model generated QA pair verification was done by a team of 14 annotators. Combined, this entire effort took around 1,170 MWH.
IV HCT-QA Synthetic Generator
Figure 5: HCT-QA Synthetic Data Generator workflow
To enable large-scale generation of HCTs with their corresponding QA pairs, we built a synthetic generator comprising two components: a HCT Data Generator (Section IV-A) and a QA Generator (Section IV-B). This generator serves as a valuable resource for the research community, facilitating the exploration of research challenges on HCTs. The generator is highly configurable, allowing it to capture nearly all common HCT and question properties defined earlier. It can also be configured to create QA pairs with multiple operators, which would be difficult to create manually at scale.
The main goal of the HCT-QA Synthetic Generator is to reduce the time and labor cost required to create HCTs and QA pairs on them. The entire process of manually extracting real-world HCTs, creating manually generated or human-verified model-generated QA pairs, and labeling the HCT and question properties takes around 4 minutes per QA pair. In contrast, creating the domain vocabulary for an entire family of synthetic HCTs, specifying the related table and question templates takes about 1 hour in total with the generation of a synthetic HCT instance and a related QA pair from that family taking only less than a second. Thus, the synthetic generator enables scaling up the generation process to several thousands HCTs and QA pairs in a few hours, while still controlling their diversity and complexity through the template design. As we will show throughout this section, the HCT-QA Synthetic Generator has the following four properties:
  1. 1.
    Scalability: It can be easily extended to cover new domains and generate thousands of HCTs and QA pairs with minimal efforts.
  2. 2.
    Guaranteed semantic correctness: Based on a template-based approach, semantic fidelity can be ensured across the constructed HCTs and their generated NL questions.
  3. 3.
    Broad coverage of question complexity: NL questions are generated using various SQL templates spanning different levels of sophistication and operator complexities.
  4. 4.
    Disambiguation: The generated NL questions have a clear interpretation in terms of what they are referring to in the table regardless of their complexity.
The synthetic HCT and QA generation process (Figure 5) begins with domain specifications (a) to create custom table templates (b) and question templates (d) from which, inspired by SQL, sample instances of relational tables $T_{REL}$ and SQL queries $Q_{SQL}$ are generated, respectively. Executing $Q_{SQL}$ on $T_{REL}$ gives ground-truth results $A$ (e). We derive HCT templates (c) from relational table templates to control how to pivot and style the same $T_{REL}$ instance into different variants of HCTs instances $T_{HCT}$. We derive NL templates (f) from SQL templates that shape the transcription of $Q_{SQL}$ queries into NL questions $Q_{NL}$. Results $A$ (e) to $(T_{REL},Q_{SQL})$ instances are used as answers to $(T_{HCT},Q_{NL})$ instances as they are identical by design. Metadata $M$ characterizing HCTs and question properties (Section III) are computed too, forming the $(T_{HCT},Q_{NL},A,M)$ benchmark quadruplets.
Figure 6: (a) Statistics of the synthetic HCTs and QAs for 77 domains. An example of a relational table $T_{REL}$ (b) and a generated $T_{HCT}$ from it (c). (d) Examples of synthetic QAs, where $Q_{SQL}$ is transformed into $Q_{NL}$ using custom templates, and executed on $T_{REL}$ to get the ground-truth answer $A$.
IV-A Synthetic HCT Data Generator
The HCT Data Generator generates HCTs in multiple formats such as CSV, HTML, PDF, and PNG, and supports various stylistic features such as row header indentation, various background colors and fonts, and line separations. Through a combination of configuration files and templates, the generator can reproduce $15$ (marked with “*” in Table II) of the $19$ HCT properties. The remaining 4 properties are not supported in this version due to their complex irregularities.
Domain Specifications (Figure  5(a)). We begin by defining a set of realistic application domains, such as import–export statistics across countries and months over several years. Each domain is associated with a domain vocabulary consisting of (1) a JSON configuration file listing attribute names and their set of possible values and (2) catalog tables listing a comprehensive set of attributes from the vocabulary, their data types, and possible value ranges. Various HCTs are then created by different combinations of attributes and value ranges. For example, in a “Food import–export” domain (Figure 6(b)), user-defined attributes such as Year: {2015,...,2023}, ImportExport Op.: {Import, Export}, Food Category: {Meat, Dairy}, and Food Item: {Chicken, Beef, Cream, Milk, Yogurt} are used to generate HCTs describing food import-export activities over years.
Types of Attributes (Figure  5(a)). We distinguish between two types of attributes, independent attributes such as Year and Import-Export for which all pairs of values are possible, e.g., (2022, Import), and hierarchical attributes such as (Country, State, City) and (Food category, Food item), where attribute values follow an inherent order and only certain combinations are meaningful. For example, combinations such as (California, Boston) are invalid. Independent attributes can appear as either row or column headers at any nesting level. In contrast, hierarchical attributes must respect logical order — parent attributes, such as Country, must precede child ones, such as City, from left to right or top to bottom within headers. Moreover, all attributes belonging to the same hierarchy must be placed together, either entirely in the row headers or entirely in the column headers, e.g., an HCT cannot have Country as a row header and City as a column header. More details are provided in [18].
Relational Table Templates (Figure 5(b)). We manually define Table templates in JSON format that give details about which attributes and values to pick for a given scenario to form a family of relational tables. For example, for the import-export scenario, the data generator uses the table template to create the relational table ($T_{REL}$) depicted in Figure 6(b). The template defines a target field, e.g., the Value numerical attribute in Figure 6(b), to contain the fine-grained value (automatically generated) for each combination over the rest of the attributes. The generator enforces uniqueness constraints on the target field to prevent ambiguity when evaluating model-predicted answers against ground-truth values. Relational tables serve two essential purposes: (1) They enable systematic generation of HCTs through pivoting operations [24, 53]; and (2) they are used to construct the ground truth answers by executing SQL queries on them without human intervention.
HCT Templates (Figure 5(c)). These templates define the final construction rules for producing HCTs instances from the domain specifications and structural combinations set in the relational table template from which they are automatically generated. They are highly configurable, for example, one can specify the depth of row and column headers, the assignment of attributes to each header, and the inclusion and type of aggregation functions and whether the aggregation occurs at a global or local level. They also capture layout specifications such as nesting balance and symmetry, as well as stylistic elements, e.g., line separations, and indentation.
As shown in Figure 6(c), an example HCT instance ($T_{HCT}$) is generated from its corresponding relational table ($T_{REL}$). In this case, attributes (Import_Export, Year) form nested column headers, while (Category, Items) define nested row headers. The column nesting is balanced and symmetric, whereas the row nesting is balanced but asymmetric. This instance includes both local and global aggregation rows and columns (average). Food Category row headers (Beverage and Dairy) are indented and their row contains implicit local average aggregation of food Items (e.g., Coffee and Tea for Beverage), while global aggregation of food Category and column aggregations are all explicit (”Average” appears in the header). These properties are automatically recorded in the metadata $M$. It is worth mentioning that with the relational table and the HCT templates, the Synthetic HCT Data Generator only requires the domain vocabulary to generate HCTs for new domains.
Statistics on Synthetic HCTs. In our benchmark and as summarized in Figure 6(a), we covered seven distinct domains and generated hundreds of HCTs per domain with a total of 4,679 synthetic HCTs. These HCTs exhibit different complexities and layouts as presented in Figure 3(c).
IV-B Synthetic Question & Answer Generator
Our synthetic QA Generator coupled with the HCT Generator, allows us to scale HCT-QA to produce tens of thousands of QA on thousands of synthetic HCTs with $100\%$ accuracy.
Figure 7: Example of a SQL query template, a basic NL template, and its instantiation on the HCT (top) for synthetic query generation.
SQL Query Templates (Figure 5(d). The QA Generator builds upon the same domain specifications and table templates described in Section IV-A. Table III lists 15 SQL query templates, each representing a distinct level of complexity. These templates are used to instantiate SQL queries with varying structures and operations from simple or multiple rows and columns selection (SQL1,2,3,6), to multiple aggregations over one row or one column (SQL4,7), when the same aggregation is already displayed in the HCT (SQL5), or grouped by any attribute levels or only by the top level one, and with or without reporting the attributes (SQL8,9,10,11). We also generate sorting and top-k queries (SQL12,13) as well as selections with operations (SQL14) and a selection conditioned on another selection (SQL15).
As an example, Figure 7 presents a SQL query template (SQL7 in Table III) to generate global aggregation queries (i.e., without sub-groups) on the relational representation $T_{REL}$ (not shown) of an HCT. The template applies one equality condition to each attribute involved in column nesting, allowing it to naturally extend to multi-level nesting scenarios. When applied to the presented HCT, these conditions trace through the structure, one level at a time, to reach a specific column (green & blue colors of the WHERE clauses correspond to the same color boxes in the HCT column headers). Additionally, the template introduces IN conditions on attributes associated with row nesting. The HCT in Figure 7 contains no row nesting, and hence a maximum of one condition is possible (i.e., on the Month column). In contrast, the HCT in Figure 6(c) contains row nesting, and the template allows putting conditions on the corresponding columns (i.e., on the Category and Item columns in $T_{REL}$). These conditions contribute to increasing the query’s complexity. Finally, the template applies one or more aggregation functions over the selected output values.
Ground Truth Generation (Figure  5(e). The relational table $T_{REL}$ and the generated SQL queries are imported into a relational DBMS for execution and the generation of the ground-truth answer $A$, thus forming the triplet ($T_{REL},Q_{SQL},A$). In Figure 6(d), we illustrate three SQL query instances generated from different SQL query templates (SQL9, SQL12, and SQL15 in Table III).
Natural Language Templates (Figures 5(d-f)). For each SQL query template, we manually define a NL template to transcribe $Q_{SQL}$ into $Q_{NL}$ specific to the HCT domain. To illustrate the need for these custom NL templates, let’s consider the SQL query template in Figure 7 that would be transcribed into a (basic) NL template as:
"What is the \$AGG1 and \$AGG2 of Values for \$Col_Level1 equals \$v1 and \$Col_level2 equals \$v2 and \$RowNest_Level1 IN <list>"
Then, a (basic) NL question instance, given some values from the HCT in Figure 7, would be:
"What is the SUM and AVG of Values for Airlines equals ‘Other Airlines’ and Direction equals ‘Arrival’ and Month IN [‘January’, ‘March’, ‘May’]"
Obviously, this basic direct transcription does not form a seamless NL question. Hence, we customize the NL templates for each of the seven domains covered in our generator (Figure 6(a)), generating more natural questions. We illustrate a few examples in Figure 6(d). Similar to the construction of synthetic HCTs, the Synthetic QA generator is domain-agnostic. It relies solely on specifying a domain by its vocabulary describing attributes (e.g. Month) and values (May, June…) in a JSON format, and the manual design of $15$ $Q_{NL}$ templates based on this vocabulary and simple rules to ensure English syntax. Each $Q_{NL}$ template transcribes in natural language one of the $15$ $Q_{SQL}$ templates shown in Table III, ensuring non-repetitive queries. The full list of $15$ $Q_{NL}$ templates for each of the $7$ domains and the rules to build them, is available in Appendix D in the extended report [18]. As each $Q_{NL}$ template corresponds to a $Q_{SQL}$ template with specific query properties, we automatically record them in the metadata $M$.
TABLE III: SQL templates of Synthetic HCT-QA. $r|c$: depth of row||column (R||C) headers of the HCT. $Attr_{C}^{*}$: all attributes of C. $Attr_{R}^{1}$: left-most attribute of R. NumNum: avg. of selected column. $Dir\stackrel{{\scriptstyle\text{rand}}}{{\sim}}\{$ASC, DESC}\}, $k\stackrel{{\scriptstyle\text{rand}}}{{\sim}}\{2,3,4,5\}$, and $Op\stackrel{{\scriptstyle\text{rand}}}{{\sim}}\{<,>\}$.
SQL templates
SQL1 SELECT Value FROM Data WHERE (AttrR = ValR)r AND (AttrC = ValC)c
SQL2 SELECT Value FROM Data WHERE (AttrR IN ListValR)r AND (AttrC = ValC)c
SQL3 SELECT Value FROM Data WHERE (AttrR = ValR)r AND (AttrC IN ListValC)c
SQL4 SELECT AGG1(Value), AGG2(Value) FROM Data WHERE (AttrR = ValR) AND (AttrC IN ListValC)
SQL5 SELECT AGG(Value) FROM Data WHERE (AttrR = ValR)r AND (AttrC IN ListValC)c
SQL6 SELECT Value FROM Data WHERE (AttrR IN ListValR)r AND (AttrC IN ListValC)c
SQL7 SELECT AGG1(Value), AGG2(Value) FROM Data WHERE (AttrR IN ListValR)r AND (AttrC = ValC)c
[…]
SQL11 SELECT $_{R}^{1}, Attr$$_{C}^{*}$, AGG1(Value) FROM Data WHERE (AttrR IN ListValR)r AND (AttrC IN ListValC)c GROUP BY $_{R}^{1}, Attr$$_{C}^{*}$
SQL12 SELECT Value FROM Data WHERE (AttrR IN ListValR)1 AND (AttrC = ValC)c ORDER BY VALUE Dir LIMIT k
SQL13 SELECT Value FROM Data WHERE (AttrR IN ListValR)1 AND (AttrC = ValC)c ORDER BY VALUE Dir
[…]
Statistics on Synthetic Questions. We generated 67,747 QA pairs on the $4,679$ synthetic HCTs across seven domains.
Scalability. Thanks to manually defined JSON templates, our Synthetic Data Generator can be augmented with minimal effort. The domain specifications can be augmented at will, adding attribute names and their list of values, to generate HCTs in any domain of interest. Also, the relational table templates allow controlling the diversity and complexity of the generated HCTs. SQL query templates can be easily configurable due to their computational nature. The corresponding NL templates require minimum effort to adapt to a new domain, using a small set of syntax rules. This template-based automation allows unlimited HCT-QA accurate data generation in minimal time without the need for costly verification.
Manual Effort. A key advantage of the synthetic data generator is that, once the tool is implemented—a process that required several hundred hours of development—adding new SQL or natural language templates for a specific domain typically takes no more than two to three hours.
V Experimental evaluation
While HCT-QA is designed to evaluate data processing engines during answering NL questions on HCTs, our evaluation focus on LLMs and VLMs. To this end, we explain our experimental setup in Section V-A and then answer seven research questions (RQs) experimentally in Section V-B.
V-A Experimental setup
@@
Models. We tested a large variety of open- and closed-weight LLMs and VLMs (Table IV). All open-weight models were downloaded from HuggingFace.
Datasets. There are 1,880 real-world HCTs from four data sources along with 9,835 QA pairs and 4,679 synthetic HCTs with 67,747 QA pairs (total of 6,559 HCTs with 77,582 QA pairs). For the results in Table IV, the text modality HCTs were represented in HTML format and the vision-text modality HCTs were images encoded in base64.
@@
Prompting Approach and Hyperparameters. We experimented with zero-shot and one-shot prompting with no clear indication that the latter improved model performance. Thus, all subsequent experiments use zero-shot prompting. We ask the model to ”only answer the question using information from the HCT ” and allow it to refuse to answer the question if it deems it unanswerable. For inference, the temperature was set to 0, and the maximum output tokens to 256. For finetuning, the learning rate was 5.0e-07 and the number of epochs was 2. The finetuning config file can be found in [18]. We post-process the model responses using a rule-based approach to extract the actual answer if one was provided. All inference and finetuning experiments were run on 8 NVIDIA A100 GPUs. Inference which required around 77k prompts took a total of 129 hours, averaging 1 hour, 1.5 hours, and 22 hours per small, medium, and cloud model, respectively. Additionally, our finetuning experiments took 30 hours.
Evaluation Metrics. We use two metrics for model accuracy: (i) F1 Score–commonly adopted in several table QA benchmarks, e.g.,  [65, 11, 71, 69, 13], and (ii) Complete Containment (CC Score)–a binary score that evaluates to 1 iff the recall is 100%, otherwise it is 0. The two metrics are complementary: F1 rewards partial correctness of the answer while CC rewards an answer only if it contains the ground truth in its entirety. For example, assume the ground truth is [“Apple”, “300”] and three distinct models output: M1 [“Apple”, “300”], M2 [“Apple”], and M3 [“Apple”, “300”, “Orange”, “10”], then their scores will be: M1 (F1= 1.0, CC= 1), M2 (F1= 0.66, CC= 0), and M3 (F1= 0.66, CC= 1).
TABLE IV: Overall performance of models on HCT-QA
@@
Model Name Model Param Size QNPC PAK US CENSUS arXiv Synthetic Tables Average
F1 CC Score F1 CC Score F1 CC Score F1 CC Score F1 CC Score F1 CC Score